OpenStudy (ineedhelp10):

If the coordinates of a square that is inscribed in a circle are (-3,6), (5,6), (-3,-2), and (5,-2) what is the area of the circle?

OpenStudy (anonymous):

If a square is inscribed in a circle, then the diagonal of the square is the diameter. So take any of the two diagonals, say (5,6) and (-3,-2) and find the lngth of this diagonal using the distance formula. So that distance is the diameter. The radius is half the diameter. Then use the formula, Area of a circle = pi(radius)^2.

OpenStudy (ineedhelp10):

@Easyaspi314 Then the diameter will be 8?

OpenStudy (anonymous):

One sec..let me see.

OpenStudy (anonymous):

I get the diagonal equal to sqrt(128) = 8sqrt(2)

OpenStudy (anonymous):

so radius is half of 8sqrt(2) = 4 sqrt(2).

OpenStudy (anonymous):

Use the distance formula to be convinced that the diameter is sqrt(128).

OpenStudy (ineedhelp10):

so you mean \[sqrt{128}\]

OpenStudy (anonymous):

yes, and sqrt 128 = sqrt 64 times sqrt 2 = 8 sqrt 2

OpenStudy (ineedhelp10):

so techinally 128 is the diameter and 64 is the radius?

OpenStudy (anonymous):

not 128, sqrt 128 is the diameter!

OpenStudy (anonymous):

sqrt 128 is the diameter! radius is half of sqrt 128 = 4 sqrt(2)

OpenStudy (anonymous):

@ineedhelp10

OpenStudy (ineedhelp10):

I dont get how you got 4 sqrt(2)

OpenStudy (anonymous):

Do you agree that the diameter is sqrt(128)?

OpenStudy (anonymous):

@ineedhelp10

OpenStudy (anonymous):

well, I assume you're not here.

OpenStudy (ineedhelp10):

Yeah I agree with that and I am here

OpenStudy (ineedhelp10):

idk how you got 4 as the radius tho?

OpenStudy (ineedhelp10):

I don get why whould you square 64 to get the radius?

OpenStudy (ineedhelp10):

@Easyaspi314

OpenStudy (ineedhelp10):

are you there?

OpenStudy (ineedhelp10):

@e.mccormick can you please explain this? the person who tried explaning it logged off

OpenStudy (e.mccormick):

So basically this, right? 

OpenStudy (ineedhelp10):

yeah, i get that part

OpenStudy (e.mccormick):

Where the diagonals cross is the center. That means from the center to the corner is a radius. Then you can use the radius formula. 

OpenStudy (e.mccormick):

radius to area formula, that is.

OpenStudy (ineedhelp10):

Okay.I understand that part too

OpenStudy (e.mccormick):

Kk. So where do you get messed up at?

OpenStudy (ineedhelp10):

So I know the diagonal is 8 right? and after that I used the Pythagorean Theorom to find the diameter which is 128, but im confused on finding the radius

OpenStudy (e.mccormick):

The diagonal is \(8\sqrt{2}\)

OpenStudy (ineedhelp10):

Thats not how they taught it to me in class

OpenStudy (e.mccormick):

Take a triangle a, b, c for sides with c as the hypotenuse: \(a^2+b^2=c^2\) But this is frm a square, so a=b. \(a^2+a^2=c^2\) \(2a^2=c^2\) \(\sqrt{2a^2}=c\) \(a\sqrt{2}=c\) Therefore since a is 8, c is \(8\sqrt{2}\)

OpenStudy (e.mccormick):

It is a rule for any 45-45-90 triangle.

OpenStudy (ineedhelp10):

I think you just confused me even more

OpenStudy (e.mccormick):

Well, you said you used the Pythagorean. I did too. I can do it with numbers and show you also.

OpenStudy (ineedhelp10):

it'll be better to show me with number plz

OpenStudy (e.mccormick):

\(8^2+8^2=c^2\) \(64+64=c^2\) now, I could say that is 128 on the left, but because I know that the square root of 64 is 8, I am going to do it this way instead: \(2\cdot 64=c^2\) Do you get that?

OpenStudy (ineedhelp10):

Why would you go back and multiply 2 by 64?

OpenStudy (e.mccormick):

I can write 128 lots of way. That is one. \(128 = 2\cdot 8\cdot 8\) \(128 = 2\cdot 2\cdot 4 \cdot 2\cdot 4 \) \(128 = 2\cdot 2\cdot 2\cdot 2 \cdot 2\cdot 2\cdot 2\) \(128 = 2^7\) All of them mean the same thing.

OpenStudy (ineedhelp10):

okay, so then square 128 is the diameter?

OpenStudy (e.mccormick):

Yes. But when I simplify that I get \(8\sqrt{2}\) Let me work with the all 2s. \(2\cdot 2\cdot 2\cdot 2 \cdot 2\cdot 2\cdot 2=c^2\) \(\sqrt{2\cdot 2\cdot 2\cdot 2 \cdot 2\cdot 2\cdot 2}=c\) Now, I can take them out in pairs because I am taking them out from under a square root. \(2\sqrt{2\cdot 2\cdot 2\cdot 2 \cdot 2\cdot \cancel{ 2\cdot 2}}=c\) \(2 \cdot 2\sqrt{2\cdot 2\cdot 2\cdot \cancel{ 2\cdot 2}}=c\) \(2 \cdot 2\cdot 2\sqrt{2\cdot \cancel{ 2\cdot 2}}=c\) Ah, and \(2 \cdot 2\cdot 2=8\) so: \(2 \cdot 2\cdot 2\sqrt{2}=c\) becomes \(8\sqrt{2}=c\) And sode c of the triangle is the diametter of the circle.

OpenStudy (ineedhelp10):

This is confusing to me because my geometry teacher didnt teach it this way to us.

OpenStudy (e.mccormick):

Hmm... Well, you have to be able to at least divide it in half. Or just use the divided form. \(\sqrt{128}\) means the same thing, so if you want it in that form it is OK. In fact, since you use \(\pi r^2\) that is not a bad form.

OpenStudy (e.mccormick):

That is all I am sayig. It is the same number. Bit of algebra there.

OpenStudy (ineedhelp10):

oh okay thank you. can you explain it this way instead please

OpenStudy (e.mccormick):

OK. so if \(D=\sqrt{128}\) and \(r=\dfrac{D}{2}\), with me there?

OpenStudy (ineedhelp10):

yeah, i get you

OpenStudy (e.mccormick):

\(r=\dfrac{\sqrt{128}}{2}\) Now it is just plugging that into the area formula.

OpenStudy (e.mccormick):

\(\pi r^2 = \pi \left(\dfrac{\sqrt{128}}{2}\right)^2\) Can you solve that?

OpenStudy (ineedhelp10):

so the radius will be \[\sqrt{64}\]

OpenStudy (ineedhelp10):

@e.mccormick

OpenStudy (e.mccormick):

Ummm... \(\sqrt{64}=8\) and \(\dfrac{\sqrt{128}}{2} = \dfrac{11.31}{2}\) which is less than 6. So they are not the same.

OpenStudy (e.mccormick):

For you to divide, it would have to be under a root on the bottom.

OpenStudy (ineedhelp10):

the radius will be 8, correct?

OpenStudy (e.mccormick):

No.

OpenStudy (e.mccormick):

You can't divide like that.

OpenStudy (ineedhelp10):

When I put \[\sqrt{128}\] in my calculator I get 11.31

OpenStudy (e.mccormick):

Yes.

OpenStudy (e.mccormick):

5.655

OpenStudy (ineedhelp10):

I didnt divide by 2 tho and I got the answer

OpenStudy (ineedhelp10):

so bottomline, the radius is 5.65?

OpenStudy (e.mccormick):

\(r\approx 5.655\) is correct. But that is not \(\sqrt{64}\). Now, does your teacher want an exact answer or an approximate?

OpenStudy (ineedhelp10):

I think the exact answer

OpenStudy (e.mccormick):

OK. Then there are two ways you can do that. Square it first, then divide, or do the division then square it. Because it is \(\pi r^2\) it has to be squared in there somewhere.

OpenStudy (ineedhelp10):

My teacher didnt divide \[\sqrt{128}\] by 2 in another problem tho. So will you need to divide the sqaure root for this type of problems each time?

OpenStudy (e.mccormick):

Yoy can write it as \(r=\dfrac{\sqrt{128}}{2}\) and be just fine. Then when you do \(\pi r^2\) the root goes away, but the bottom gets squared too!

OpenStudy (e.mccormick):

Also, in the other problem, you may have had the radius. This time you have the diameter. So you have to change the diameter to the radius, which is where the over 2 comes from.

OpenStudy (ineedhelp10):

My teacher said the answer was 32pi for the area, is that correct?

OpenStudy (e.mccormick):

For this? Yes.

OpenStudy (ineedhelp10):

I dont get that when plugging it in my calculator though

OpenStudy (e.mccormick):

\(r=\dfrac{\sqrt{128}}{2}\) means: \(A =\pi r^2 = \pi \left(\dfrac{\sqrt{128}}{2}\right)^2\) Right? No need for a calculator.

OpenStudy (ineedhelp10):

Oh okay I think I understnad. But how come in other type of problems you leave it in terms or pi?